OpenStudy (anonymous):

one plant has a mass of 15 grams and increases its mass by 3 grams each day. Another plant has a mass of 7 grams and it increases its mass at the rate of 5 grams each day. On what day will they have the same mass?

OpenStudy (anonymous):

15 + 3x = 7 + 5x 8 = 2x x = 4 so, 4 days

OpenStudy (anonymous):

15,18,21,.... 7,12,17,.... nth day they will have same mass This is an arithmetic sequence. nth term of 15,18,21,..... will be 12 +3n [ a = 15, d = 3, tn =a +(n-1)d = 15 +(n-1) 3 = 15 +3n -3 = 12 +3n ] nth term of 7,12,17,.... 2 +5n [ a = 7, d = 5, tn =a +(n-1)d = 7+(n-1)5 = 7 +5n -5 = 2 +5n ] Equate both 12 +3n = 2 +5n Solve for n 10 = 2n 10/2 = n 5 = n 5th day they will have equal mass 15,18,21,........24,27 7,12,17,.........22,27

OpenStudy (anonymous):

I don't know how you got 5 honestly. It is 4, do it with the math i gave or write them out.

OpenStudy (anonymous):

Oh I see what you did wrong .. when you got n=5, you forget that the answer of an arithmetic series is (n-1).